The annual output and prices of a 3-good economy are shown in the table below. Assume Year 1 is the base year. 


Quantity of Ice CreamPrice of Ice CreamQuantity of ShampooPrice of ShampooQuantity of Peanut ButterPrice of Peanut Butter
Year 13$51$43$3
Year 25$92$62$4


What was nominal GDP in year 1?$ 

What was real GDP in year 1?$ 

What was nominal GDP in year 2?$ 

What was real GDP in year 2? $
Answer #1

Answer
Nominal GDP =sum of ( prices of the year* quantities of the year)
Nominal GDP in year 1=3*5+1*4+3*3=28
real GDP =sum of (the price of the base year * quantity of the year)
real GDP in year 1=3*5+1*4+3*3=28
nominal GDP in year 2=5*9+2*6+2*4=65
real GDP in year 2=5*5+2*4+2*3=39